| Operation Summary for January 11, 2012 - January 13, 2012 |
|
Wednesday January 11 The day shift began with Operations monitoring the Linac, Booster, and Main Injector (MI), and with the NuMI and MiniBooNE experiments taking beam. During the day shift, Booster personnel began a study at 9:27 AM. At 10:46 AM, the MI-10 kickers tripped off; they reset. Operations established beam to the new Meson FTBF experiment T-922 at 11 AM. At 11:23 AM, a Linac RF station (LRF3) tripped off two times in 13 minutes; it reset. Muon Rings personnel began a high intensity study at 1:31 PM. At 2:41 PM, the MI-10 kickers tripped off for the first of two times in about 45 minutes; after the second trip it took many attempts before the kickers actually reset. During the evening shift, LRF3 tripped off for the first of five times in seven hours; it reset each time. The Muon Rings study ended at 4:15 PM. A Recycler person began a study at 4:31 PM. A MI RF station (MIRF9) tripped off and would not reset; Operators bypassed the station and brought MIRF6 on in its place. Operations disabled T-922 beam at 6:38 PM. The MI-10 kickers tripped off at 7:05 PM and would not reset; a MI kicker person investigated and reset the kickers locally. Operators conducted a MI tune at 9:23 PM. At 10:46 PM, a Booster vacuum pump (IPRF5) tripped off and caused BRF3 and BRF4 to also trip off; everything reset. Thursday January 12 The midnight shift began with Operations sending beam to the NuMI and MiniBooNE. During the midnight shift, at 12:13 AM, BRF17 and BRF18 tripped off for the first of seven times in three hours; they reset. Operations loaded protons into the Recycler at 1:45 AM. At 1:57 AM, three MI kicker (KPS1A, B, & C) high voltage power supplies tripped off; they reset. At 5:30 AM, the NuMI beam permit tripped off due to an incomplete RF gap indication; it reset. During the day shift, at 8:11 AM, a Booster vacuum pump (IPRF22) tripped off and caused BRF11 and BRF12 to also trip; everything reset. A Recycler person began a study and a Booster person began tuning. Operations established beam to T-922 at 10:53 AM. A Muon Rings person began a two-hours study at 11:37 AM. At noon, the MI-10 kickers tripped off for the first of two times in 40 minutes; they reset. Operations reported at 12:53 PM that the Neutron Therapy Facility (NTF) conducted a one-hour calibration run. Operators performed a 750 KeV Line tune at 1:15 PM. The MI-10 kickers tripped off at 1:23 PM; they reset. During the evening shift, the Recycler studies ended at 5:58 PM. Operators loaded protons into the Recycler at 6:44 PM. At 6:49 PM, LRF3 tripped off twice in four minutes; it reset. At 7:09 PM, BRF17 and BRF18 tripped off on a power supply overload indication; they reset. Operations disabled beam to T-922 at 10:03 PM. At 10:11 PM, the Booster West Anode Power Supply tripped off and would not reset; Operators investigated and managed to get the anode supply on, but then it tripped off again. Further investigation found that BRF12 was the cause; Operators bypassed the station. BRF17 also tripped off, but with help from a RF person, Operators brought the station online. A Linac Klystron RF station (KRF5) tripped off at 11:12 PM; it reset. Friday January 13 The midnight shift began with Operations sending beam to the NuMI and MiniBooNE. During the midnight shift, Operators made an adjustment to the Booster transition time at 1:40 AM due to BRF12 being down. At 3:35 AM, MIRF16 tripped off; it reset. At 5:30 AM, LRF4 tripped off; it reset. At 7:39 AM, NTF began preparing to treat one patient later in the morning. Operations Coordinator (OpsCo) Plans for Today and the Weekend The OpsCo approved of the plans by Booster and MI personnel to conduct studies. The OpsCo also reported that on 1/18/12, there would be an eight-hour access in the Booster tunnel for solid-state modulator and power amplifier upgrades. Other machines may also conduct maintenance. |
|
More Information Comments and Suggestions |
| last modified 01/13/2012 email Fermilab |
| Security, Privacy, Legal | Use of Cookies | |